OpenStudy (adunb8):

can anyone explain to be the differential equations: first order (exact equations) ? for example (3x^2y^2-4xy)y' + 2xy^3-2y^2=0

OpenStudy (lgbasallote):

first...turn y' into dy/dx \[\implies (3x^2 y^2 - 4xy)\frac{dy}{dx} + 2xy^3 - 2y^2 = 0\] now...multiply dx to all sides \[\implies (3x^2 y^2 - 4xy) dy + (2xy^2 - 2y^2)dx = 0\] following so far?

OpenStudy (adunb8):

yea

OpenStudy (adunb8):

then what happens?

OpenStudy (lgbasallote):

now you see if it's exact

OpenStudy (adunb8):

how?

OpenStudy (lgbasallote):

\[\frac{\partial M}{\partial y} = \frac{\partial N}{\partial x}\] for \[Mdx + Ndy = 0\] in this case.. your M = (2xy^3 - 2y^2) and N = 3x^2 y^2- 4xy so \[\frac{\partial M}{\partial y} = 6xy^2 - 4y\] \[\frac{\partial N}{\partial x} = 6xy^2 - 4y\] so it's exact got it?

OpenStudy (adunb8):

i see so what happens after that? do i just say it is exact because i think my teacher wants be to show that it is a implicit solution or something

OpenStudy (lgbasallote):

you solve it now... take the \[\int Mdx\] so in this case.. \[\implies \int (2xy^3 - 2y^2)dx\] \[\implies\int 2xy^3 dx - \int 2y^2 dx\] \[\implies x^2y^3 - 2y^2 x + g(y)\] (that g(y) is kind of like the arbitrary constant c) now derive it WITH RESPECT TO Y \[\implies 3xy^2 y^2 - 4y^2 x + g^\prime(y)\] still following?

OpenStudy (adunb8):

yes !!! omg!

OpenStudy (lgbasallote):

uhh wait that should be \[\implies 3x^2 y^2 - 4xy + g^\prime (y)\] lots of typos

OpenStudy (adunb8):

oh okay lol

OpenStudy (lgbasallote):

now equate this to N so.. \[\implies 3x^2 y^2 - 4xy + g^\prime (y) = 3x^2 y^2 - 4xy\] cancel the cancellable... \[\implies g^\prime (y) = 0\] now integrate both sides \[\int g^\prime (y) = \int 0\] \[\implies g(y) = 0\] substitute this into the g(y) written earlier... \[\implies x^2 y^3 - 2y^2 x + g(y)\] if you substitute \[\implies x^2 y ^3 - 2y^2 x + 0\] \[\implies x^2 y^3 - 2y^2 x\] this is now the final answer

OpenStudy (adunb8):

oh i see! so integrate the Mdx and then equal to N and see if they are equal wow this was a great lesson! thank you so much!

OpenStudy (unklerhaukus):

great work Lgba

OpenStudy (lgbasallote):

welcome

OpenStudy (lgbasallote):

and thanks @UnkleRhaukus

OpenStudy (adunb8):

wow i think i should learn more from you im actually taking differential equation & linear algebra this semester and i think it would really help out if you keep updating my questions =) Thank You so much!